if $10,000 is invested in the stock at the start of 2015 how much will it be come by year end 2018

are 2015 10% return year 2016 - 5% return year 2017 12% return and year 2018 15% return
Answer #1

investment value = $10,000

2015:

return for the year = 10%

so gain = 10,000 x 10% = 1000

total amount at the end of year 2015 = 10,000 + 1000 = 11,000

2016:

amount at the beginning = $11,000

return for the year = -5%(loss)

loss = 11,000 x 5% = $550

total amount at the end of 2016 = 11,000 - 550 = $10,450

2017:

amount at the beginning = 10,450

return for the year = 12%

gain = 10450 x 12% = $1254

total amount the end of year = 10450 + 1254 = $11,704

2018:

amount at the beginning = $11,704

return for the year = 15%

gain = 11,704 x 15% = 1755.6

so total amount at the end of year 2018 = 11,704 + 1755.6 = $13,459.6